# T-cell receptor specificity landscape revealed through de novo peptide design

### v1

@@ III Results
Next, we examined the therapeutic T-cell responses to all single point mutants of the HLA-A*02:01-restricted p53-derived neoepitope p53R175H, HMTEVVRHC, measured in ref. [5]. This T-cell therapy approach used a designed bi-specific antibody that binds both to the TCR-CD3 complex on T-cells and to the variants of the p53R175H-MHC complex on tumor cells (Table S1). A strong binding to an epitope activates the engaged T-cell, and the extent of such activity was measured by the production of IFN-$\gamma$. The resulting HERMES-fixed predicted peptide energies (eq. 1) show a Spearman $\rho=46\%$ with the experimentally measured IFN-$\gamma$ production (Figs. 3D, S2).
Our analysis shows that, despite not being explicitly trained for this purpose, HERMES peptide energy effectively predicts T-cell responses to neoepitopes presented by different MHC-I alleles. Moreover, HERMES predictions are not limited to TCR-pMHC interactions but also extend to antibody-pMHC interactions, commonly leveraged in T-cell therapies for their higher achievable affinities. The HERMES predictions on T-cell activities perform as well or better than other computational methods including BLOSUM62 [47], ProteinMPNN [30], and TCRdock [17]; see Table S4 for a detailed benchmark of different approaches, including different HERMES models.
@@
Vaccination protocols aim to elicit robust immune responses that induce lasting immune memory in individuals against viral infections and cancer. Peptide-based cancer vaccines represent a promising therapeutic strategy to mount an effective T-cell response, offering the potential to be highly personalized to an individual’s unique tumor antigens and T-cell repertoire [55, 7]. However, the clinical success of cancer vaccines has been limited by factors such as low peptide immunogenicity and insufficient uptake by antigen-presenting cells [56].
Recent approaches have employed deep mutational scanning (DMS) experiments to identify more immunogenic cancer neoantigens [36]. However, these experiments are often limited to scanning single amino acid mutations from the wild-type (native) peptide due to the combinatorial explosion when exploring multiple mutations.
@@
Here, we present a pipeline for designing peptides with enhanced TCR-MHC reactivity by leveraging HERMES’s ability in predicting T-cell response to peptides presented by MHC-I molecules. Our structure-based approach starts with an existing template structure–experimentally or computationally resolved–of a TCR-MHC in complex with at least one peptide variant, which we refer to as the wild-type peptide. We use two methods to sample peptides, with different degrees of complexity; see Fig. 1C for a schematic description of the two pipelines.
@@
In the basic pipeline (HERMES-fixed), we sequentially mask the atoms of each amino acid along the peptide, one at a time. Using the HERMES model, we predict the probability of each amino acid type for a given residue, based on the local structural environment within 10 Å  of the masked residue. Repeating this procedure for every amino acid along the peptide yields a position weight matrix (PWM) that represents the probabilities of different amino acids at each peptide position. Peptides are then sampled by drawing from this PWM. It should be noted that HERMES-fixed retains the structural fingerprints of the original peptide, as it does not account for local structural changes induced by amino acid substitutions; see Fig. 1C and SI for more details.
@@
To incorporate structural flexibility and relaxation following amino acid substitutions, we introduce the design pipeline of HERMES-relaxed, using simulated annealing and Markov chain Monte Carlo (MCMC) sampling. We begin with the template TCR-MHC structure and a completely random peptide sequence, which is packed in silico within the structure using PyRosetta’s packing functionality, followed by the Relax protocol [28]. During relaxation, both side-chain and backbone atoms of the peptide are allowed to move, while only the side-chain atoms of the MHC and TCR chains within 10 Å of the peptide are flexible. With this relaxed structure, we apply the HERMES-fixed method to generate a PWM and sample a new peptide, which is then packed and relaxed to form a new structure. We use MCMC to determine whether to accept the newly sampled peptide by comparing its HERMES peptide energy (favorability), evaluated in its relaxed conformation, to that of the original peptide (eq. 1). We then iteratively perform this procedure, incrementally reducing the MCMC temperature at each step, until the results converge; see Figures 1C, and S3, and SI for more details.
@@
The peptides designed by both approaches are then filtered by TCRdock [17] or AlphaFold3 [29], using the Predicted Alignment Error (PAE) between TCR and pMHC interfaces, to assure folding and binding of the resulting TCR-pMHC structures. For each system, we define a specific PAE threshold close to the TCRDock PAE evaluated for the system’s TCR-MHC in complex with the native (wild-type) peptide, which is known a priori to activate T-cells. As a result, the PAE thresholds vary slightly across different systems; see SI for details.
@@
We tested the accuracy of our peptide design pipeline in three systems, with the structural templates taken from (i) the cancer-testis antigen NY-ESO-1 in complex with the TCR 1G4 and HLA-A*02:01, (ii) a peptide derived from the Epstein-Barr virus (EBV) in complex with the TK3 TCR and the HLA-B*35:01, and (iii) the immunotherapeutic target MAGE-A3 in complex with an engineered TCR and HLA-A*01:01. For brevity, we refer to these three systems as NY-ESO, EBV and MAGE, respectively; See Table S5 for information on the specific templates used in each case. It should be noted that as we progressed from NY-ESO to EBV and MAGE, we chose to explore the sequence space farther from the wild-type peptides, leading to more challenging designs.
@@
We validated our designs using Jurkat cells expressing a specific TCR and an endogenous NFAT-eGFP reporter to indicate T-cell activation in response to different peptides. We tested the TCR-MHC specificity in all systems–NY-ESO, MAGE, EBV–under 96 different conditions, including the de novo designed peptides, positive controls (wild-type peptides from the template structures), and an unstimulated “no peptide” control. Peptides were presented on artificial APCs (aAPCs), expressing the specified HLA in each system. For NY-ESO designs, we determined the the percentage of GFP-positive cells through both flow cytometry and fluorescence microscopy. Given the consistency between the two experimental approaches (Fig. S4) we relied on fluorescence microscopy only to measure GFP levels induced by different peptides for MAGE and EBV designs; see SI for experimental details.
@@
We used two structural templates for the NY-ESO system: one containing the original 9-amino acid peptide SLLMWITQC, and another with the more immunogenic C9V substitution [46, 34]. Our de novo designs differed from the closest wild-type peptide by 2 to 7 amino acids, with a TCRDock PAE $\leq 5.5$. For this system only, we selected 35 negative designs (designed by HERMES, but with PAE $>5.5$) for experimental testing: only one of these was a false negative, which interestingly, had a desirable AF3 PAE score (as low as the wild-type peptides); see Figs. S5, S6. For the 58 positive designed peptides (PAE  $\leq\leavevmode\nobreak\ 5.5$), we achieved an overall design accuracy of 50%, with GFP levels significantly higher than the negative control; see SI for details. The accuracy of predictions decreases as the sequence divergence of the de novo designs from the wild-type peptides increases (Fig. 4A, S6). HERMES-fixed achieved higher design accuracy of 70% within its smaller designed sequence subspace of up to 5 substitutions from the wild-types, whereas HERMES-relaxed explored sequences further from the wild-type peptides (4-7 substitutions) but with a reduced accuracy of 24% (Fig. 4A). The main difference between the true and false positive sequences was at position 8–from strongly preferring glutamine within the true positives, to having similar preferences among glutamine, glycine, arginine, or methionine within the false positives (Fig. 4A).
@@
For the EBV system, we used two structural templates: one with the 11-amino acid peptide HPVGEADYFEY—commonly referred to as HPVG—derived from the viral latent nuclear protein EBNA-1, and the other with an epitope from a wide-spread viral variant, with a glutamine at position 5 [52]. Unlike NY-ESO, the HPVG peptide has a more flexible conformation with a helical turn within the MHC binding pocket. Our de novo designs differed from the closest wild-type peptide by 3 to 9 amino acids, with a TCRDock PAE  $\leq\leavevmode\nobreak\ 5.1$. Across the 93 de novo designs, we achieved an overall accuracy of 14%, associated with the peptides that induced significant GFP levels in T-cells. The accuracy of predictions decreases with increasing the sequence divergence from the wild-types (Fig. 4B, S7). Notably, we achieved 40% accuracy among the designs within 3-5 amino acid distance of one of the wild-type templates, while none of the designs with larger than 5 amino acid distance were successful. Lastly, we observe a strong preference for glutamic acid at position 10 of the successful designs relative to the false positives (Fig. 4B).
@@
For the MAGE system, we used two structural templates: one with the 9-amino acid MAGE-A3 peptide EVDPIGHLY and another with the Titin-derived self-peptide ESDPIVAQY, which is expressed in cardiomyocytes [53]. The Titin-derived peptide is known to have triggered an off-target response of the affinity-enhanced MAGE-A3-specific engineered TCR that resulted in severe cardiac toxicity and deaths of two patients [57]. Our de novo designs differed from the closest wild-type peptide by 3 to 8 amino acids, with a TCRDock PAE  $\leq\leavevmode\nobreak\ 5.2$ (Fig. 4C). From the resulting 93 de novo designs, 20% activated T-cells with significant levels of GFP expression. Similar to the other systems, the accuracy of predictions decreases with increasing the sequence divergence from the wild-types (Fig. 4C). We achieved 30% accuracy among the designs with 3-5 to amino acid distance, while none of the designs with larger than 5 amino acid distance from the wild-types were successful.
@@
Fig. 4C shows differences in amino acid composition between successful and failed designs. Alanine and glycine scanning experiments have previously suggested that MAGE-A3 peptide variants with an E-DPI- - -Y motif generally activate the T-cells in this system [58]. Consistent with these findings, the fraction of T-cell activating variants in our designs increases as their sequences more closely resemble this motif (Fig. S8).
@@
Although no direct experimental comparisons were performed, computational evaluations using TCRdock PAE and presentation score by MHC-I alleles indicate that HERMES designs outperform or match those generated by ProteinMPNN [30], the corresponding MHC position weight matrix [59, 60, 61, 54], or the BLOSUM substitution matrix [47]; see SI and Figs. S9, S10, and S11 for details. A notable exception is in the MAGE system, where ProteinMPNN designs generally have more favorable PAE scores, albeit with a lower proportion of designs passing the antigen presentation filter by the MHC molecule (Fig. S11).
In all the three systems we used the PAE values from TCRdock and AF3 as filtering criteria to select designs for experimental validation (see SI). The PAE values reported by the structure prediction algorithms have been shown to be noisy indicators of protein folding fidelity and design quality [62]. While larger experimental libraries are needed to draw definitive conclusions, it appears that PAEs from TCRdock and AF3 have some complementary predictive strengths (Figs. S5, S12), and therefore, combining their information could lead to a more robust design decisions (SI).
Overall, HERMES demonstrates a potential for designing highly immunogenic peptides up to five amino acid substitutions away from the wild-type–a task that would otherwise require exploring a vast sequence space of roughly $10^{8}$ to $10^{10}$ possibilities for a 9-residue to an 11-residue peptide (Fig. 4A-C). This capability makes HERMES a powerful platform for the de novo design of immunogenic peptides and neoantigens, offering a new computational path for the development of peptide-based T-cell vaccines against viruses and cancer.
@@
Our computational peptide design framework helps address this limitation, at least for the TCR-MHC complexes with a known structural template. Specifically, for a given TCR-MHC pair, we can us HERMES-fixed model to generate a peptide position weight matrix (PWM), representing the ensemble of peptides presented by the MHC and recognized by the TCR. We use the entropy of this PWM as a proxy for TCR degeneracy.
@@
Because our analyses focus on interactions involving typical TCR and MHC molecules, our entropy-based degeneracy estimates are smaller than those from previous theoretical studies, which often relied on experimental observations from autoimmune T-cells or limited synthetic peptide libraries [65]. Moreover, the HERMES-fixed model generates PWMs based on peptides whose backbone conformations match those of the wild-type template, likely underestimating the true degeneracy of TCR-MHC complexes. Although HERMES-relaxed can partially address this issue, its designs are only reliable within roughly five amino acids of the native peptides (Fig. 4), preventing a full exploration of the possible peptide poses for a given TCR-MHC complex. Nonetheless, our estimate can serve as informative lower bounds on typical MHC-I and TCR degeneracies.
IV Discussion

### v2

@@ Results
Next, we examined the therapeutic T-cell responses to all single-point mutants of the HLA-A*02:01-restricted p53-derived neoepitope p53R175H, HMTEVVRHC, measured in ref. [5]. This T-cell therapy approach used a designed bi-specific Fab antibody that binds both to the TCR-CD3 complex on T-cells and to the variants of the p53R175H-MHC complex on tumor cells (Table S1). A strong binding to an epitope activates the engaged T-cell, and the extent of such activity was measured by the production of IFN-$\gamma$. The resulting HERMES-fixed predicted peptide energies (eq. 1) show a Spearman $\rho=0.46$ with the experimentally measured IFN-$\gamma$ production (Figs. 3D, S3). ProteinMPNN achieves a slightly higher correlation (Spearman $\rho=0.55$), but this improvement over HERMES is not statistically significant (Fisher’s z-test p-value >0.05>0.05); see Table 1. Lastly, we found that replacing the crystal structure with the AF3 model substantially degrades HERMES and ProteinMPNN performances (Fig. S4, Tables 1, S3), as the AF3 model consistently mis-docks the bispecific antibody on the pMHC complex (Table S4 and Fig. S5).
Across the limited number of systems examined, our analyses indicate that–when a suitable structural template is available–HERMES can estimate epitope mutational effects at the pMHC-TCR interface and, in the H2-scDb case, at a Fab-pMHC interface. In this benchmark, ProteinMPNN performs comparably to HERMES, indicating that inverse-folding algorithms not explicitly trained for this task may nonetheless capture the impact of peptide mutations on T-cell responses. However, definitive conclusions about the generalizability of these methods to neoepitopes will require larger datasets with broader mutational coverage across peptide variants. Importantly, predictive performance for both approaches depends on access to a high-quality structure (experimental or computational), which limits applicability in systems without suitable structural models (Fig. S4); see SI and Tables 1 and S3 for a detailed benchmark of different approaches, including different HERMES models.
@@
TCR-engineered T-cell therapies are designed to target defined neoantigens, but specificity remains a central challenge. Engineered high-affinity TCRs can cross-react with untested self-peptides, causing severe off-target toxicity; in a MAGE-A3 directed trial, recognition of a Titin-derived peptide led to fatal cardiac events [59]. Therefore, developing approaches to predict the distribution of peptides recognized by engineered TCRs could enable more stringent preclinical screening and help mitigate off-target risk in patients.
@@
To map the peptide-recognition landscape of TCRs, we present a structure-guided pipeline for designing immunogenic peptides predicted to elicit strong TCR recognition when presented by MHC class I. Leveraging HERMES’s ability to predict T-cell responses to peptides presented by MHC-I, our approach begins with an existing template structure–experimentally or computationally resolved—of a TCR-MHC in complex with at least one peptide variant, which we refer to as the wild-type peptide. We then generate candidate peptides using two strategies, with different degrees of complexity; see Fig. 1C for a schematic description of the two pipelines.
@@
In the basic pipeline (HERMES-fixed), we sequentially mask the atoms of each amino acid along the peptide, one at a time. Using the HERMES model, we predict the probability of each amino acid type for a given residue, based on the local structural environment within 10 Å  of the masked residue. Repeating this procedure for every amino acid along the peptide yields a position weight matrix (PWM) that represents the probabilities of different amino acids at each peptide position. Peptides are then sampled by drawing from this PWM. It should be noted that HERMES-fixed retains the structural fingerprints of the original peptide, as it does not account for local structural changes induced by amino acid substitutions; see Fig. 1C and S6A, and SI for more details.
@@
To incorporate structural flexibility and relaxation following amino acid substitutions, we introduce the design pipeline of HERMES-relaxed, using simulated annealing and Markov chain Monte Carlo (MCMC) sampling. We begin with the template TCR-MHC structure and a completely random peptide sequence, which is packed in silico within the structure using PyRosetta’s packing functionality, followed by the Relax protocol [29]. During relaxation, both side-chain and backbone atoms of the peptide are allowed to move, while only the side-chain atoms of the MHC and TCR chains within 10 Å of the peptide are flexible. With this relaxed structure, we apply the HERMES-fixed method to generate a PWM and sample a new peptide, which is then packed and relaxed to form a new structure. We use MCMC to determine whether to accept the newly sampled peptide by comparing its HERMES peptide energy (favorability), evaluated in its relaxed conformation, to that of the original peptide (eq. 1). We then iteratively perform this procedure, incrementally reducing the MCMC temperature at each step, until the results converge; see Fig.s 1C, S6B-C, and SI for more details.
@@
The peptides designed by both approaches are then filtered by TCRdock [18] or AlphaFold3 [30], using the Predicted Alignment Error (PAE) between TCR and pMHC interfaces, to assure folding and binding of the resulting TCR-pMHC structures. For each system, we define a specific PAE threshold close to the TCRDock PAE of the system’s TCR-MHC in complex with the native (wild-type) peptide, which is known a priori to activate T-cells. As a result, the PAE thresholds vary slightly across different systems; see SI for details.
@@
We tested the accuracy of our peptide design pipeline in three systems, with the structural templates taken from (i) the cancer-testis antigen NY-ESO-1 in complex with the TCR 1G4 and HLA-A*02:01, (ii) a peptide derived from the Epstein-Barr virus (EBV) in complex with the TK3 TCR and the HLA-B*35:01, and (iii) the immuno-therapeutic target MAGE-A3 in complex with an engineered TCR and HLA-A*01:01. For brevity, we refer to these three systems as NY-ESO, EBV and MAGE, respectively; See Table S5 for information on the structure templates used in each case. These systems were selected because each one has high-resolution crystal structures of the same TCR-MHC scaffold bound to at least two distinct peptide variants, and that they span a therapeutically relevant spectrum that includes both engineered (for MAGE and NY-ESO) and natural (for EBV) TCRs directed against viral as well as tumor antigens. It should be noted that as we progressed from NY-ESO to EBV and MAGE, we chose to explore the sequence space farther from the wild-type peptides, leading to more challenging designs.
@@
We validated our designs using Jurkat cells expressing a specific TCR and an endogenous NFAT-eGFP reporter to indicate T-cell activation in response to different peptides. We tested the TCR-MHC specificity in all systems–NY-ESO, MAGE, EBV–under 96 different conditions, including the de novo designed peptides, positive controls (wild-type peptides from the template structures), and an unstimulated “no peptide” control. Peptides were presented on artificial APCs (aAPCs), expressing the specified HLA in each system. For NY-ESO designs, we determined the the percentage of GFP-positive cells through both flow cytometry and fluorescence microscopy. Given the consistency between the two experimental approaches (Fig. S7) we relied on fluorescence microscopy only to measure GFP levels induced by different peptides for MAGE and EBV designs; see SI for experimental details and Dataset S2 for data on these measurements.
@@
We used two structural templates for the NY-ESO system: one containing the original 9-amino acid peptide SLLMWITQC, and another with the more immunogenic C9V substitution [51, 38]. Our de novo designs differed from the closest wild-type peptide by 2 to 7 amino acids, with a TCRDock PAE $\leq 5.5$. For this system only, we selected 35 negative designs (designed by HERMES, but with PAE >5.5>5.5) for experimental testing: only one of these was a false negative, which interestingly, had a desirable AF3 PAE score (as low as the wild-type peptides); see Figs. S8, S9. For the 58 positive designed peptides (PAE  $\leq\leavevmode\nobreak\ 5.5$), we achieved an overall design accuracy of 50%, with GFP levels significantly higher than the negative control; see SI for details. The accuracy of predictions decreases as the sequence divergence of the de novo designs from the wild-type peptides increases (Fig. 4A, S9). HERMES-fixed achieved higher design accuracy of 70% within its smaller designed sequence subspace of up to 5 substitutions from the wild-types, whereas HERMES-relaxed explored sequences further from the wild-type peptides (4-7 substitutions) but with a reduced accuracy of 24% (Fig. 4A). The main difference between the true and false positive sequences was at position 8–from strongly preferring glutamine within the true positives, to having similar preferences among glutamine, glycine, arginine, or methionine within the false positives (Fig. 4A).
@@
For the EBV system, we used two structural templates: one with the 11-amino acid peptide HPVGEADYFEY—commonly referred to as HPVG—derived from the viral latent nuclear protein EBNA-1, and the other with an epitope from a wide-spread viral variant, with a glutamine at position 5 [56]. Unlike NY-ESO, the HPVG peptide has a more flexible conformation with a helical turn within the MHC binding pocket. Our de novo designs differed from the closest wild-type peptide by 3 to 9 amino acids, with a TCRDock PAE  $\leq\leavevmode\nobreak\ 5.1$. Across the 93 de novo designs, we achieved an overall accuracy of 14%, associated with the peptides that induced significant GFP levels in T-cells. The accuracy of predictions decreases with increasing the sequence divergence from the wild-types (Fig. 4B, S10). Notably, we achieved 40% accuracy among the designs within 3-5 amino acid distance of one of the wild-type templates, while none of the designs with larger than 5 amino acid distance were successful. Lastly, we observe a strong preference for glutamic acid at position 10 of the successful designs relative to the false positives (Fig. 4B).
@@
For the MAGE system, we used two structural templates: one with the 9-amino acid MAGE-A3 peptide EVDPIGHLY and another with the Titin-derived self-peptide ESDPIVAQY, which is expressed in cardiomyocytes [57]. Our de novo designs differed from the closest wild-type peptide by 3 to 8 amino acids, with a TCRDock PAE  $\leq\leavevmode\nobreak\ 5.2$ (Fig. 4C). From the resulting 93 de novo designs, 20% activated T-cells with significant levels of GFP expression. Similar to the other systems, the accuracy of predictions decreases with increasing the sequence divergence from the wild-types (Fig. 4C). We achieved 30% accuracy among the designs with 3-5 to amino acid distance, while none of the designs with larger than 5 amino acid distance from the wild-types were successful.
@@
Fig. 4C shows differences in amino acid composition between successful and failed designs. Alanine and glycine scanning experiments have previously suggested that MAGE-A3 peptide variants with an E-DPI- - -Y motif activate T-cells in this system [60]. Consistent with these findings, the fraction of T-cell activating variants in our designs increases as their sequences more closely resemble this motif (Fig. S11).
@@
To gauge the impact of using AF3 structures as design templates, we reran the HERMES-fixed protocol with AF3 predicted structures (with and without template guidance), and obtained peptide PWMs that were almost identical to those derived from using crystal structures (SI and Fig. S12). An interesting exception is the MAGE system, where using the AF3 template introduced a strong preference for glutamic acid at positing 1 of the peptide (Fig. S12C), mirroring the pattern observed among the designed peptides that induced strong T-cell response (Fig. 4D).
Although no direct experimental comparisons were performed, computational evaluations using TCRdock PAE and presentation score by MHC-I alleles indicate that HERMES designs outperform or match those generated by ProteinMPNN [31], the corresponding MHC position weight matrix [61, 62, 63, 58], or the BLOSUM substitution matrix [40] (SI and Figs. S13, S14, S15). A notable exception is in the MAGE system, where ProteinMPNN designs generally have more favorable PAE scores, albeit with a lower proportion of designs passing the antigen presentation filter by the MHC molecule (Fig. S15).
In all the three systems we used the PAE values from TCRdock and AF3 as filtering criteria to select designs for experimental validation (see SI). The PAE values reported by the structure prediction algorithms have been shown to be noisy indicators of protein folding fidelity and design quality [64]. While larger experimental libraries are needed to draw definitive conclusions, it appears that PAEs from TCRdock and AF3 have some complementary predictive strengths (Figs. S8, S16), and therefore, combining their information could lead to a more robust design decisions (SI). Both of these scores correlate only weakly with HERMES (Figs. S8, S13, S14, S15), so they can provide complementary design filters. Furthermore, designed peptides with higher (more favorable) HERMES energies more often activated T-cells (Fig. S16), suggesting that a more restrictive sampling (e.g., with a lower MCMC temperature for HERMES-relaxed protocol) could yield a higher success rate, albeit at the expense of peptide diversity.
Overall, HERMES demonstrates a potential for designing highly immunogenic peptides up to five amino acid substitutions away from the wild-type–a task that would otherwise require exploring a vast sequence space of roughly $10^{8}$ to $10^{10}$ possibilities for a 9-residue to an 11-residue peptide (Fig. 4A-C). This capability raises an important question: to what extent do these substantial changes in sequence alter the binding geometry of the peptide inside the MHC-TCR cleft? To this end, we used AF3 to model the TCR-pMHC interfaces for the experimentally validated immunogenic designs from the three systems. Although we see slight variation in both peptide and TCR backbone conformations, overall, we found that our designed immunogenic peptides adopt conformations very close to their wild‑type templates, even when those designs were generated with HERMES‑relaxed (Fig. 4E). We attribute this structural convergence to two potential factors: (i) residual limitations in HERMES peptide‑backbone sampling strategy, and (ii) an inherent AF3 bias toward conformations it encountered during training. Given that TCRs can cross-react with peptides of markedly different conformational poses [65], developing methods that can better explore peptide conformational space within the TCR-MHC groove remains a compelling direction for future research.
@@
Our computational peptide design framework helps address this limitation, at least for the TCR-MHC complexes with known structural templates. Specifically, for a given TCR-MHC pair, we can use HERMES-fixed protocol to generate a peptide position weight matrix (PWM), representing the ensemble of peptides presented by the MHC and recognized by the TCR. We use the entropy of this PWM as a proxy for TCR degeneracy.
@@
As an alternative to MHC Motif Atlas [58], we explored the ability of HERMES-fixed to model MHC degeneracy. When structural coverage is extensive (roughly >30>30 structures of distinct peptides for a given MHC allele), the motifs produced by HERMES align closely with those in the MHC Motif Atlas (SI, Fig. S17, and Table S7). With sparser structural coverage, HERMES underestimates the peptide diversity, likely because backbone conformations are insufficiently sampled. Given the imbalanced distribution of structures across MHC alleles, we therefore continue to use MHC Motif Atlas [58] as our reference for computing MHC motif specificity.
Because our analyses focus on interactions involving typical TCR and MHC molecules, our entropy-based degeneracy estimates are smaller than those from previous theoretical studies, which often relied on experimental observations from autoimmune T-cells or limited synthetic peptide libraries [68]. Moreover, the HERMES-fixed model generates PWMs based on peptides whose backbone conformations match those of the wild-type template, likely underestimating the true degeneracy of TCR-MHC complexes (Fig. 5). Although HERMES-relaxed can partially address this issue, its designs are only reliable within roughly five amino acids of the native peptides with similar backbone conformations (Fig. 4), preventing a full exploration of the possible peptide poses for a given TCR-MHC complex. Nonetheless, our estimates can serve as informative lower bounds on typical MHC-I and TCR degeneracies.
Discussion